# Interpretability-Guided Bi-objective Optimization: Aligning Accuracy and Explainability

### v1

@@ III-D Gradient-Based Optimization via Projection Mapping
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$: for every $\lambda\in(0,1)$ work
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$: just some of $\lambda\in(0,1)$ guarantees both
@@
$\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=\begin{cases}\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2},\quad 0\leq g_{1}\cdot g_{2},\lambda\in(0,1)\\ \lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2},g_{1}\cdot g_{2}<0,\lambda\in(0,1)\end{cases}$ (13)
@@
Geometric Intuition: The function $\mathcal{P}$ provides a complete parameterization of the simultaneous descent convex combinations $C_{\text{SD}}=\{\mathbf{v}=\alpha g_{1}+\beta g_{2}:\alpha,\beta>0,\,\mathbf{v}\cdot\mathbf{g}_{1}>0\text{ and }\mathbf{v}\cdot\mathbf{g}_{2}>0\}$:
@@
Unified Theoretical Guarantee (Complete Characterization): From Theorems A-A3 and A-A4:
@@
Convergence Guarantees: Using Theorems A-A3 and A-A4, we establish:
@@
    Robustness: The method tolerates gradient estimation errors (Theorem A-B1).
@@
This projection-based approach transforms the bi-objective optimization in IGBO from a heuristic balancing act to a principled procedure with full theoretical guarantees, enabling stable training with explicit trade-off control.
@@ IV-A Computational Complexity Analysis
Training with IGBO introduces additional computational overhead compared to standard training, primarily due to the computation of Temporal Integrated Gradients (TIG) and the interpretability loss. We analyze the time and memory complexity for both standard training and IGBO.
@@
    Oracle Query: finding $K$ points which satisfies the mentioned conditions $O(K\cdot G)$ .
@@
    $\textbf{Time}=O(\text{MAX}\{M\cdot T\cdot F,G,T^{2}\cdot d,d^{2}\}),$
    $\text{ since }|E|\in O(d^{2})$
@@
    • The Oracle finding each point which satisfies the mentioned conditions: $G\rightarrow O(\mathcal{G})$.
@@
  • IGBO (Parallel TIG computation + Oracle query): $O(M\cdot T\cdot h+Q)$ (for computing all $M$ points simultaneously) Additionally, with time complexity trade-off we have:
    • $\textbf{Memory}=O(M^{a}\cdot T\cdot h+Q)$
@@
Note on Geometric Update Complexity: The geometric update fucntion (Eqs. (13)) adds only $O(p)$ operations in sequential processing and $O(1)$ in parallel processing, where $p$ is the number of parameters. This is negligible compared to the $O(F)$, confirming that the geometric approach does not introduce significant computational burden beyond the interpretability loss computation.
V Experimental Design
This section outlines the comprehensive experimental plan to evaluate the IGBO framework. All experiments will be conducted following this protocol to ensure rigorous, reproducible validation.
V-A Research Objectives
Our experimental campaign aims to address four core objectives:
  • Objective 1: Constraint Enforcement Efficacy
    • Quantify how effectively IGBO enforces DAG-specified feature importance hierarchies
    • Measure the accuracy cost of enforcing interpretability constraints
  • Objective 2: Oracle Performance Analysis
    • Evaluate whether the Optimal Path Oracle generates more in-distribution integration paths compared to linear interpolation
    • Assess the impact of oracle-guided paths on attribution stability
  • Objective 3: Parameter Trade-off Characterization
    • Systematically analyze the effects of parameters $K$ (oracle points) and $M$ (total integration points)
    • Map the trade-off space between computational cost and path quality
  • Objective 4: Robustness Assessment
    • Test IGBO’s sensitivity to imperfect domain knowledge (noisy DAGs)
    • Evaluate robustness to hyperparameter variations ($\epsilon$, $\delta$, $\beta$)
V-B Dataset Preparation Plan
We will prepare and utilize three datasets:
  • Synthetic Hierarchical Dataset (SHD)
    • Generation Method: Create time-series data with known causal structure $X_{1}\rightarrow X_{2}\rightarrow X_{3}\rightarrow X_{4}$
    • Data Size: 10,000 sequences of length $T=50$ with $d=4$ features
    • Ground Truth: Precisely known feature importance hierarchy for validation
    • Split: 70% train, 15% validation, 15% test
  • MIMIC-III Clinical Dataset
    • Task: In-hospital mortality prediction within first 48 hours
    • Preprocessing: Extract 24-hour multivariate time-series, handle missing values with forward-fill, normalize features
    • DAG Construction: Collaborate with clinicians to define $G_{\text{clinical}}$: Vital signs $\succ$ Lab values $\succ$ Demographics
    • Expected Size: $\sim$15,000 patient stays after filtering
  • Financial Time-Series Dataset
    • Source: Yahoo Finance/Kaggle for S&P 500 constituents
    • Task: Next-day volatility classification (high/low)
    • Features: Technical indicators, price data, volume metrics
    • DAG: Define $G_{\text{finance}}$: Price-based $\succ$ Macroeconomic $\succ$ Sentiment
V-C Baseline Implementation Plan
We will implement and compare the following methods:
  • Primary Baselines:
    1. 1.
      Standard Model (STD): Base LSTM/Transformer trained only with task loss
    2. 2.
      Weighted Sum Regularization (WSR): $\mathcal{L}+\lambda\mathcal{H}$ with $\lambda\in\{0.001,0.01,0.1,1.0,10.0\}$
    3. 3.
      Post-hoc Fine-tuning (PHF): Two-stage: STD training followed by $\mathcal{H}$ fine-tuning
    4. 4.
      Linear Path IGBO (LP-IGBO): IGBO with $K=0$ (ablation study)
  • Implementation Details:
    • All methods use identical architecture: 2-layer LSTM, 128 hidden units, dropout=0.2
    • Adam optimizer, learning rate $5\times 10^{-4}$, batch size 32
    • Early stopping based on validation loss with patience=20
    • 5 random seeds per configuration for statistical significance
V-D Evaluation Metrics to Report
For each experiment, we will compute and report:
  • Primary Task Metrics:
    • Classification: Accuracy, F1-Score, AUROC (with confidence intervals)
    • Regression: RMSE, MAE, $R^{2}$
    • Relative performance: $\Delta_{\text{Acc}}=\frac{\text{Acc}_{\text{STD}}-\text{Acc}_{\text{Method}}}{\text{Acc}_{\text{STD}}}$
  • Interpretability Metrics:
    • DAG Satisfaction Rate (DSR): $\frac{1}{|E|}\sum\mathbb{I}(\hat{d}_{u,v}\in[\epsilon,\delta])$
    • Constraint Violation Magnitude (CVM): Average deviation from interval boundaries
    • Feature Importance Rank Correlation: Spearman correlation with ground truth (for SHD)
  • Oracle-Specific Metrics:
    • OOD Score: $\frac{1}{M}\sum_{j=1}^{M}D(\mathbf{q}_{j})$ where $D$ is pre-trained discriminator
    • Path Length Ratio: $\frac{\text{Oracle path length}}{\text{Linear path length}}$
    • Attribution Consistency: Std. dev. of $H_{k}$ scores across random seeds
    • Computational Cost: Training time per epoch, memory usage, inference time
V-E Parameter Studies Design
  • Study 1: Oracle Configuration ($K$, $M$)
    • Grid search: $K\in\{0,1,3,5,7\}$, $M\in\{10,20,50\}$
    • For each $(K,M)$: Train IGBO to convergence, record all metrics
    • Analyze: DSR vs. Computation time, OOD Score vs. $K$
  • Study 2: DAG Margin Sensitivity
    • Vary $\epsilon\in\{0.05,0.1,0.2\}$, $\delta\in\{0.3,0.5,1.0\}$
    • Fixed $K=3$, $M=20$
    • Measure: DSR, CVM, and accuracy trade-offs
  • Study 3: Robustness to Noisy Priors
    • Corruption levels: 0%, 10%, 20%, 30% of edges randomly reversed
    • Evaluate on both corrupted and uncorrupted edges separately
    • Assess recovery of correct constraints despite noise
  • Study 4: Generator Weight ($\beta$)
    • $\beta\in\{0.1,0.5,1.0,2.0,5.0\}$ in $\mathcal{L}_{G}=\mathcal{L}_{\text{path}}+\beta\mathcal{L}_{\text{valid}}$
    • Impact on path quality and training stability
V-F Experimental Procedure
  • Phase 0: Infrastructure Setup
    1. 1.
      Implement all models in PyTorch with modular design
    2. 2.
      Create data loaders for all three datasets
    3. 3.
      Implement metrics computation and logging utilities
    4. 4.
      Set up experiment tracking (Weights & Biases)
  • Phase 1: Baseline Establishment
    1. 1.
      Train Standard model on all datasets (5 seeds each)
    2. 2.
      Train WSR with all $\lambda$ values, identify Pareto frontier
    3. 3.
      Train PHF with different fine-tuning durations
    4. 4.
      Establish performance baselines
  • Phase 2: IGBO Core Evaluation
    1. 1.
      Pre-train discriminator $D$ for each dataset (50 epochs)
    2. 2.
      For each $(K,M)$ configuration in Study 1:
      • Train oracle generator $G$ (20 epochs warm-up)
      • Train full IGBO model
      • Record all metrics at convergence
    3. 3.
      Repeat for 5 random seeds per configuration
  • Phase 3: Sensitivity Analysis
    1. 1.
      Execute Study 2 (margin sensitivity)
    2. 2.
      Execute Study 3 (noisy DAG robustness)
    3. 3.
      Execute Study 4 (generator weight)
  • Phase 4: Analysis and Visualization
    1. 1.
      Statistical analysis: paired t-tests with Bonferroni correction
    2. 2.
      Generate all planned visualizations:
      • Pareto frontier plots (accuracy vs. DSR)
      • Trade-off surfaces ($K$, $M$, computation, DSR)
      • Feature importance heatmaps for case studies
      • t-SNE visualizations of integration paths
    3. 3.
      Compile results into comprehensive tables
V-G Validation Criteria
We will consider IGBO successful if:
  • Primary Success Criteria:
    • Achieves DSR ¿ 80% on test sets while maintaining ¡ 5% relative accuracy drop vs. STD
    • Oracle paths show significantly higher OOD Score than linear paths ($p<0.01$)
    • Attribution Consistency improves by ¿ 20% vs. linear paths
  • Secondary Success Criteria:
    • Parameters $K$ and $M$ show predictable trade-offs
    • Robust to moderate DAG noise (¡ 20% corruption)
    • Computationally feasible (training time ¡ 2× STD)
V-H Computational Resources and Timeline
  • Hardware:
    • 4× NVIDIA V100 GPUs (32GB memory each)
    • 64 CPU cores, 256GB RAM
    • Total estimated GPU hours: 2,000 hours
  • Software:
    • PyTorch 2.0, Python 3.9
    • Weights & Biases for experiment tracking
    • Docker containers for reproducibility
  • Timeline:
    • Months 1-2: Infrastructure and baseline implementation
    • Months 3-4: Core IGBO experiments (Phase 2)
    • Months 5: Sensitivity analyses (Phase 3)
    • Month 6: Analysis, visualization, and paper writing
V-I Reproducibility Measures
To ensure reproducibility:
  • All code will be publicly released on GitHub
  • Docker images with full environment specifications
  • Configuration files for every experiment
  • Pre-trained models and evaluation scripts
  • Detailed lab notebook documenting all experimental decisions
This experimental design provides a complete roadmap for evaluating IGBO. The structured approach ensures comprehensive assessment of all framework components while maintaining scientific rigor and reproducibility.
VI Conclusion and Future Work
We have introduced Interpretability-Guided Bi-objective Optimization (IGBO), a novel framework for training machine learning models to adhere to domain-specific interpretability constraints. By formalizing these constraints via a Directed Acyclic Graph (DAG) acting on a differentiable feature importance function $H_{k}(\mathbf{X},\theta)$, and by deriving a geometrically motivated update rule that dynamically decomposes task and interpretability gradients, IGBO provides a principled method to navigate the accuracy-interpretability trade-off. Theoretical analysis guarantees convergence to stationary points with clear semantic meanings: joint optima, task-only optima, or Pareto-optimal trade-off points.
The empirical framework we have outlined is designed to validate IGBO’s effectiveness in enforcing structured priors with minimal accuracy loss, outperforming naive regularization baselines. The ability to explicitly encode relational domain knowledge (“feature A should be more important than feature B”) and have the learning algorithm respect it during training represents a significant step towards more transparent, trustworthy, and alignable models for high-stakes applications.
Limitations and Future Work:
  • DAG Specification: The current framework requires the DAG and intervals $\epsilon,\delta]$ to be specified by an expert. Future work could explore learning or refining this graph from data, or from interactions with a human-in-the-loop.
Appendix A Proofs
This appendix contains the complete proofs for the theoretical claims made in Section III-D.
A-A Proofs for Projection Mapping Properties
A-A1 Preliminary Definitions

### v2

@@ III-E Gradient-Based Optimization via Projection Mapping
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$: for every $\lambda\in(0,1)$ works
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$: just some of $\lambda\in(0,1)$ guarantee both
@@
$\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=\begin{cases}\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2},\quad 0\leq\mathbf{g}_{1}\cdot\mathbf{g}_{2}\\ \lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2},\quad\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0\end{cases}$ (13)
@@
Geometric Intuition: The function $\mathcal{P}$ provides a complete parameterization of the simultaneous descent convex combinations $C_{\text{SD}}=\{\mathbf{v}=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}:\alpha,\beta>0,\,\mathbf{v}\cdot\mathbf{g}_{1}>0\text{ and }\mathbf{v}\cdot\mathbf{g}_{2}>0\}$:
@@
Unified Theoretical Guarantee (Complete Characterization): From Theorems V-B3 and V-B4:
@@
Convergence Guarantees: Using Theorems V-B3 and V-B4, we establish:
@@
    Robustness: The method tolerates gradient estimation errors (Theorem V-C1).
@@
This projection-based approach transforms the bi-objective optimization in IGBO from a heuristic balancing act to a principled procedure with full theoretical guarantees, enabling stable training with explicit trade-off control. Algorithm 1 provides the complete implementation.
@@ IV-A Computational Complexity Analysis
Training with IGBO introduces additional computational overhead compared to standard training, primarily due to the computation of TIG and the interpretability loss. We analyze the time and memory complexity for both standard training and IGBO.
@@
    Oracle Query: finding $K$ points that satisfy the mentioned conditions $O(K\cdot G)$ .
@@
    $\textbf{Time}=O(\text{MAX}\{M\cdot T\cdot F,K\cdot G,T^{2}\cdot d,d^{2}\}),$
    $\text{ since in simple DAGs }|E|\in O(d^{2})$
@@
    • The Oracle finding each point that satisfies the mentioned conditions: $G\rightarrow O(\mathcal{G})$.
@@
  • IGBO (Parallel TIG computation + Oracle query): $O(M\cdot T\cdot h+\mathcal{Q})$ (for computing all $M$ points simultaneously) Additionally, with time complexity trade-off we have:
    • $\textbf{Memory}=O(M^{a}\cdot T\cdot h+\mathcal{Q})$
@@
Note on Geometric Update Complexity: The geometric update function (Eqs. (13)) adds only $O(p)$ operations in sequential processing and $O(1)$ in parallel processing, where $p$ is the number of parameters. This is negligible compared to the $O(F)$, confirming that the geometric approach does not introduce significant computational burden beyond the interpretability loss computation.
V Proofs
This section contains the complete proofs for the theoretical claims made in Section III.
V-A Graph Construction via CLT
V-A1 Preliminaries and notation
Let $H_{k}:\mathcal{X}\to[0,1]$ be measurable (and differentiable in $\theta$ where needed). For a single batch $B=\{\mathbf{X}^{(i)}\}_{i=1}^{n}$ sampled i.i.d. from $\mathcal{D}$ define
$\mu_{k}:=\mathbb{E}_{\mathbf{X}\sim\mathcal{D}}[H_{k}(\mathbf{X},\theta)],\qquad s_{k}(B):=\frac{1}{n}\sum_{i=1}^{n}H_{k}(\mathbf{X}^{(i)},\theta).$
For a pair $(u,v)$ set
$D^{(i)}:=H_{u}(\mathbf{X}^{(i)},\theta)-H_{v}(\mathbf{X}^{(i)},\theta),$
$d_{u,v}(B):=s_{u}-s_{v}=\frac{1}{n}\sum_{i=1}^{n}D^{(i)}.$
V-A2 Lemma 2 (CLT for bounded importance scores)
Claim. If $\{\mathbf{X}^{(i)}\}_{i=1}^{n}$ are i.i.d. from $\mathcal{D}$ and $\mathrm{Var}(H_{k}(\mathbf{X},\theta))<\infty$, then
$s_{k}\;\xrightarrow{d}\;\mathcal{N}\!\left(\mu_{k},\;\frac{\mathrm{Var}(H_{k})}{n}\right)\quad(n\to\infty).$
Proof.
Apply the classical CLT to the i.i.d. bounded variables $H_{k}(\mathbf{X}^{(i)},\theta)$. Since $\mathbb{E}[H_{k}]=\mu_{k}$ and $\mathrm{Var}(H_{k})<\infty$,
$\sqrt{n}\,(s_{k}-\mu_{k})\Rightarrow\mathcal{N}(0,\mathrm{Var}(H_{k})).$
Rescaling yields the stated asymptotic distribution for $s_{k}$. Boundedness ensures finite moments and that any Gaussian mass outside $[0,1]$ vanishes as $n\to\infty$. ∎
V-A3 Lemma 3 (Mean and variance of shared-batch differences)
Claim. With $D^{(i)}$ as above,
$\mathbb{E}[d_{u,v}(B)]=\mu_{u}-\mu_{v},$
$\mathrm{Var}\big(d_{u,v}(B)\big)=\frac{1}{n}\Big(\mathrm{Var}(H_{u})+\mathrm{Var}(H_{v})-2\mathrm{Cov}(H_{u},H_{v})\Big).$
Proof.
Linearity of expectation gives $\mathbb{E}[d_{u,v}]=\mathbb{E}[H_{u}]-\mathbb{E}[H_{v}]=\mu_{u}-\mu_{v}$. Independence across samples implies
$\mathrm{Var}\!\Big(\frac{1}{n}\sum_{i=1}^{n}D^{(i)}\Big)=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathrm{Var}(D^{(i)})=\frac{1}{n}\mathrm{Var}(D^{(1)}).$
Expanding $\mathrm{Var}(H_{u}-H_{v})=\mathrm{Var}(H_{u})+\mathrm{Var}(H_{v})-2\mathrm{Cov}(H_{u},H_{v})$ yields the result. ∎
V-A4 Theorem 4 (Probability threshold equivalence)
Claim. Suppose $d_{u,v}(B)$ is approximated by
$d_{u,v}(B)\approx\mathcal{N}\!\big(\mu_{uv},\sigma_{uv}^{2}\big),\quad\mu_{uv}:=\mu_{u}-\mu_{v},$
$\sigma_{uv}^{2}:=\frac{\mathrm{Var}(H_{u})+\mathrm{Var}(H_{v})-2\mathrm{Cov}(H_{u},H_{v})}{n}.$
Then for any $\alpha>0.5$,
$P(s_{u}>s_{v})>\alpha\quad\Longleftrightarrow\quad\mu_{uv}>z_{\alpha}\,\sigma_{uv},$
and in particular
$P(s_{u}>s_{v})>0.5\quad\Longleftrightarrow\quad\mu_{u}>\mu_{v}.$
Proof.
Under the normal approximation,
$P(s_{u}>s_{v})=P(d_{u,v}>0)=1-\Phi\!\Big(\frac{0-\mu_{uv}}{\sigma_{uv}}\Big)=\Phi\!\Big(\frac{\mu_{uv}}{\sigma_{uv}}\Big).$
Thus $P(s_{u}>s_{v})>\alpha$ iff $\Phi(\mu_{uv}/\sigma_{uv})>\alpha$, i.e. $\mu_{uv}/\sigma_{uv}>z_{\alpha}$. Setting $\alpha=0.5$ gives $z_{0.5}=0$ and the stated equivalence. ∎
V-A5 Theorem 5 (Transitivity under margin/variance conditions)
Claim. Let $\alpha>0.5$. If
$P(s_{u}>s_{v})>\alpha\quad\text{and}\quad P(s_{v}>s_{w})>\alpha,$
and the corresponding margin inequalities hold
$\mu_{u}-\mu_{v}\geq z_{\alpha}\,\sigma_{uv},\qquad\mu_{v}-\mu_{w}\geq z_{\alpha}\,\sigma_{vw},$
then, using the triangle inequality for $L^{2}$ norms,
$\sigma_{uw}\leq\sigma_{uv}+\sigma_{vw},$
and consequently $P(s_{u}>s_{w})>\alpha$.
Proof.
From the margin inequalities,
$\mu_{u}-\mu_{w}=(\mu_{u}-\mu_{v})+(\mu_{v}-\mu_{w})\geq z_{\alpha}(\sigma_{uv}+\sigma_{vw}).$
Define $X:=H_{u}-H_{v}$ and $Y:=H_{v}-H_{w}$; then $H_{u}-H_{w}=X+Y$. Interpreting standard deviations as $L^{2}$ norms, $\sigma_{uv}=\|X\|_{2}$, $\sigma_{vw}=\|Y\|_{2}$, and $\sigma_{uw}=\|X+Y\|_{2}$. By Minkowski (triangle inequality in $L^{2}$),
$\sigma_{uw}=\|X+Y\|_{2}\leq\|X\|_{2}+\|Y\|_{2}=\sigma_{uv}+\sigma_{vw}.$
Combining the two displayed inequalities yields
$\frac{\mu_{u}-\mu_{w}}{\sigma_{uw}}\geq\frac{z_{\alpha}(\sigma_{uv}+\sigma_{vw})}{\sigma_{uw}}\geq z_{\alpha},$
hence $P(s_{u}>s_{w})=\Phi\big((\mu_{u}-\mu_{w})/\sigma_{uw}\big)\geq\Phi(z_{\alpha})=\alpha$. ∎
V-A6 Corollary 3 (Acyclicity of the induced graph)
Claim. The orientation rule $u\to v$ iff $P(s_{u}>s_{v})>0.5$ induces a strict total order consistent with $\{\mu_{k}\}$ and therefore a DAG. For $\alpha>0.5$, if all observed pairwise decisions satisfy the margin/variance conditions of Theorem V-A5, the induced graph is acyclic.
Proof.
For $\alpha=0.5$, Theorem V-A4 implies $u\to v$ iff $\mu_{u}>\mu_{v}$. Ordering nodes by decreasing $\mu_{k}$ yields a strict total order; orienting edges according to this order cannot produce directed cycles.
For $\alpha>0.5$, assume every observed directed edge satisfies the margin condition $\mu_{i}-\mu_{j}\geq z_{\alpha}\sigma_{ij}$. By repeated application of Theorem V-A5 along any directed path, the composed margin accumulates while the composed standard deviation is bounded by the sum of pairwise standard deviations; thus every directed path implies the corresponding direct pairwise inequality and cannot close into a directed cycle without violating the margin conditions. Hence the graph is acyclic. ∎
V-A7 Remarks on finite-sample corrections and estimation
When $n$ is moderate and some $\mu_{k}$ lie near 0 or $1$, the truncated-normal nature of $s_{k}$ can bias the Gaussian plug-in probability. Practical remedies include computing $P(s_{u}>s_{v})$ under a bivariate truncated-normal model, using bootstrap/permutation tests for $s_{u}-s_{v}$, or conservatively inflating $\sigma_{uv}$. These corrections vanish asymptotically.
V-A8 Estimator guidance
When multiple independent batches $B^{(1)},\dots,B^{(N)}$ are available, prefer the between-batch empirical variance of the batch means
$\widehat{\mathrm{Var}}(s_{k})=\frac{1}{N-1}\sum_{i=1}^{N}\big(s_{k}^{(i)}-\bar{s}_{k}\big)^{2}$
for uncertainty quantification. If only a single batch is available, estimate $\mathrm{Var}(H_{k})$ with the unbiased sample variance (divide by n1n-1) and obtain $\widehat{\mathrm{Var}}(s_{k})=\widehat{\mathrm{Var}}(H_{k})/n$. Use bootstrap or permutation methods when boundary effects or small $n$ make the CLT approximation questionable.
These lemmas, theorems, and remarks provide the formal backbone for the construction described in Section III-D; refer to them when invoking probabilistic claims or when justifying acyclicity and transitivity properties of the induced interpretability DAG.
V-B Projection Mapping Properties
V-B1 Preliminary Definitions